OpenStudy (anonymous):

What is the simplified form of sin(x + p)? a. cos x b. sin x c. –sin x d. –cos x

OpenStudy (amoodarya):



OpenStudy (anonymous):

so c would be my answer

OpenStudy (amoodarya):

yes understand ? if not look at second solution sin(x+y)=sinx cos y +cos x siny so sin(p+x)= sin pi cos x +cos pi sinx = 0 +(-1)sinx